OpenStudy (anonymous):

find all solutions for Sin^2x - 2sinx - 3 = 0 on the interval [0 , 2pi )

sam (.sam.):

You can factor it or use quadratic equation for that, what's your thought?

OpenStudy (anonymous):

can u show me by doing on Draw board ?

sam (.sam.):

Factor it like this, \[(\sin(x)-3)(\sin(x)+1)=0\] Then set each bracket equals to zero and solve for "x"

sam (.sam.):